A) How many of the following are found in 1500 g of benzene (C6H6)?

kg C6H6

mol C6H6

lb-mole C6H6

mole (g-mole) C

mol H

g C

g H

molecules of C6H6

B) A mixture of methanol and methyl acetate contains 45.0 wt% methanol.

Determine the g-moles of methanol in 300.0 kg of the mixture.

The flow rate of methyl acetate in the mixture is to be 100.0 lb-mole/h. What is the mixture flowrate in lbm/h?
Answer #1

Given 1500g C6H6 (Molecular weight = 78)

kg  C6H6 = 1500/1000 = 1.5kg

mol  C6H6 = 1500/78 = 19.23 gmol

lbmol  C6H6 = 19.23*0.0022 = 0.0424 lbmol

moles of C = 6*moles of  C6H6 = 6* 19.23 = 115.38 gmol

moles of H = 6*moles of  C6H6 = 6* 19.23 = 115.38 gmol

g of C = 115.38 gmol *12 = 1384.56g

g of H = 115.38 gmol *1 = 115.38g

molecules of C6H6 = 19.23 * 6.023*1023 = 1.1583*1025 molecules

Pls post one question as per guidelines.